OpenStudy (anonymous):

Solve the absolute value equation or indicate that the equation has no solution. |x+6|=9 B.{-15;3} C. {-3,15} D. zero with line through it

OpenStudy (anonymous):

@uri please help

OpenStudy (anonymous):

the answer is B: [-15, 3]

OpenStudy (anonymous):

Thank you @s.math

OpenStudy (anonymous):

Firstly we shall try to write the absolute value equation as an inequation

OpenStudy (anonymous):

-9<= x+6 <=9 considering -9<= x+6 we get x>=-15 and considering x+6<=9 we get x<=3 the intersection of both these gives the region -15<=x<=3